Determine the pH during the titration of 50.0 mL of 0.02 M HCl with 0.02 M NaOH after the addition of 35.0 mL and 65.0 mL of titrant.
Answer #1

Number of milimoles of Hel in 50m2 of solution = 50 ml x 0.02M = 0.054*0.02 moit = 0.001 mol = B 1. mmol Numben af milimolesTherefore, pH of the solution after addition 35 ml of NaOH is 2.45. Now, the pH of the solution after addition of 65.0 m2 ofi pH = -log [HTJ = - dog [0.384x16]. =-{ log(0.384) + log (16) } =-(-0-415-111 = 11:41 Thus the pH of the solution after ad